# Cohort Discovery: A Survey on LLM-Assisted Clinical Trial Recruitment

### v1

@@ 9 Conclusion
The ability of generative LLMs to generalize comes with a trade-off between reasoning and consistency. This work is the first to survey the rapidly evolving role of generative LLMs in clinical trial recruitment. We provide a modern perspective on the challenges in adopting LLMs to clinical research and point out promising directions for future work.
@@ 10 Limitations
This survey focuses on the role of advances in NLP in the critical domain of clinical trial recruitment. Given that this is a rapidly evolving field, we have made our best effort to include a comprehensive view of available resources and methods. It is possible that more sophisticated methods using the latest technology already exist (e.g., in the form of proprietary products), but are not yet made public.
Appendix A Trial-Patient Matching Systems
We show all the classical systems in Table 2 and the LLM-based systems in Table LABEL:tab:systems_overview_llm, by their task, method, data source characteristics and limitations.
Appendix B Systematic Reviews versus Surveys
A systematic review is a common practice in the medical research community with standardized reporting guidelines. The Preferred Reporting Items for Systematic Reviews and Meta-Analyses (PRISMA) Vrabel (2015) is the most widely used guideline. Just as finding patients given a trial eligibility criteria, this guideline requires strict inclusion and exclusion criteria to be applied to scientific literature to be included in a systematic review.
In contrast, the guidelines for writing surveys in our community, for instance, as outlined for authors by ACM Computing Surveys , TACL and ARR aim to draw perspectives on an evolving topic of interest.
@@ Appendix B Systematic Reviews versus Surveys
Criterion-level prediction Hamer et al. (2023)
1-shot prompt to LLM for criteria level prediction with explanation
Availability: Private (synthetic patient profile)
Size: 10
Mode: Short text
Type: Specific
Availability: Mixed (clinicalTrials.gov, EudraCT)
Size: 146 clinical trials
Type: Specific
Additional Annotation: Expert review
Majority of criterion-level errors due to incorrect reasoning (91%)
Unlu et al. (2024)
RAG-based
Availability: Private
Size: 2,276
Mode: Tables; text
Availability: Public∗∗
Size: 1
Type: Specific
RAG pipeline not published. Evaluated on a single trial.
Wornow et al. (2025)
Vector database of patient notes
Prompt (criteria, top-k patient notes) pair to LLM
Availability: Public (2018 N2C2: Cohort Selection) Evaluated on one trial
Doesn’t consider exclusion criteria
67% incorrect decision despite correct reasoning.
@@
Rybinski et al. (2024)
@@
Availability: Public (Koopman and Zuccon (2016), TREC CT 2021 & 2022)
Size: 58; 75; 50
Mode: Short text
Additional annotations: 3,621; 26,149; 26,581 trial-patient matches
Availability: Public∗∗
Size: 375,580

### v2

@@ 10 Conclusion
The task of clinical trial recruitment, that matches patients to a clinical trial via its participation eligibility criteria, benefits from knowledge aggregation and reasoning abilities of LLMs. In this survey, we critically examine the evolving role of LLM technologies in clinical research. We analyze the main components in a clinical trial recruitment process and provide a modern perspective on the challenges in adopting LLMs to clinical research, such as the size of the benchmarks, the dimensions of evaluation and data sensitivity. We hope that this serves as a valuable resource for future research.
@@ 11 Limitations
This survey focuses on the role of advances in NLP in the critical domain of clinical trial recruitment. Given that this is a rapidly evolving field, we have made our best effort to include a comprehensive view of available resources and methods. It is possible that more sophisticated methods using the latest technology already exist (e.g., in the form of proprietary products), but are not yet made public or are only available as abstracts, as is common in some medical communities, for example, the Annual Meeting of the American Society of Clinical Oncology (ASCO).
Appendix A Systematic Reviews versus Surveys
A systematic review is a common practice in the medical research community, with standardized reporting guidelines Moher et al. (2009); Page et al. (2021). This guideline has a checklist of required items in the title, abstract, introduction, methods, results, discussion and other information that every study needs to fulfil. These studies are bound by strict inclusion and exclusion criteria applied to the scientific literature considered for any assessment. This rigor is necessary in evidence-based analysis of a specific research question. A survey is more flexible and provides a board coverage of the topic being discussed. For instance, the guidelines for writing surveys in our community as outlined by ACM Computing Surveys , TACL and ARR aim to draw perspectives on an evolving topic of interest.
@@ Appendix A Systematic Reviews versus Surveys
Appendix B Reported Error Types
In the following we illustrate the the error distributions by percentages as reported by previous studies.
Errors reported in Jin et al. (2024) supplementary:
  • Incorrect reasoning: 30.7%
  • Lack of medical knowledge: 15.4%
  • Ambiguous label definition: 26.9%
  • other errors: 26.9% (self-conflicting)
Errors reported in Hamer et al. (2023):
  • Incorrect reading: 6.5%
  • Insufficient knowledge: 2.2%
  • Incorrect reasoning: 91.3%
Errors reported in Nievas et al. (2024):
  • Lack of knowledge  55%
  • Implicit criteria  15%
  • Wrong reasoning
  • Accurate reasoning, wrong decision
  • Lack of restraint when expert opinion is needed
  • Negated criteria error
Appendix C Trial-Patient Matching Systems
Table 3 gives an overview of all the classical systems in covered in this survey by their direction of approach, method, source data characteristics and limitations. Table LABEL:tab:systems_overview_llm covers all the LLM-based systems discussed in this survey by their task, method, data source characteristics and limitations. We notice the shift in the availability of patient sources towards public data in the LLM-based approaches. We also notice a shift towards trial ranking due to publicly available data.
@@ Appendix C Trial-Patient Matching Systems
Criterion-level prediction Hamer et al. (2023)
1-shot prompt to LLM for criteria level prediction with explanation
Availability: Private (synthetic patient profile)
Size: 10
Mode: Short text
Type: Specific
Availability: Mixed (clinicalTrials.gov, EudraCT)
Size: 146 clinical trials
Type: Specific
Additional Annotation: Expert review
Majority of criterion-level errors due to incorrect reasoning (91%)
Unlu et al. (2024)
RAG-based
Availability: Private
Size: 2,276
Mode: Tables; text
Availability: Public∗∗
Size: 1
Type: Specific
RAG pipeline not published. Evaluated on a single trial.
Ferber et al. (2024)
Criterion-level boolean eligibility labels
Availability: Private
Size: 51
Mode: EHR
Availability: Public∗∗
Size: 105,600
Type: Specific
Use of proprietary model limits evaluation on synthetic patient data.
Beattie et al. (2024)
Vector database of patient notes. Prompt (criteria, top-k patient notes) pair to LLM
Availability: Public (2018 N2C2: Cohort Selection) Some evaluation is on test subset (40 of 182).
Expert guidance requires manual expertise for every criterion.
Wornow et al. (2025)
Vector database of patient notes. Prompt (criteria, top-k patient notes) pair to LLM
Availability: Public (2018 N2C2: Cohort Selection) Doesn’t consider exclusion criteria
67% incorrect decision despite correct reasoning.
Saeidi (2025)
Prompt LLM with a fine-tuned BERT-based concept embeddings from patient and criteria to predict binary eligibility labels.
Availability: Public (2018 N2C2: Cohort Selection) LLM reasoning capabilities not explored.
@@
Cosine similarity between patient and trial embeddings obtained using GPT by Richmond and Deshpande (2023) and using Sentence Transformer by Lahiri et al. (2023). Availability: Public (TREC CT 2023) Very low MAP of 0.02 Richmond and Deshpande (2023) and close to zero for Lahiri et al. (2023)
Kusa et al. (2023b)
Sentence query formulation using GPT-3.5. Query enrichment using Kusa et al. (2023a). Trial-level prediction of re-ranked pairs using GPT-3.5.
Availability: Public (TREC CT 2023) Zero-shot LLM prompts provide marginal improvement over neural-rerankers
Peikos (2023)
Utilize GPT3.5 to obtain trial-level labels for final re-ranking on top of lexical and neural re-rankers
Availability: Public (TREC CT 2023) Query processing is template-based and excludes negative information.
Ferber et al. (2024)
No-SQL query formulation from patient EHR using GPT-4o for initial retrieval, followed by vector embedding re-ranking.
Availability: Private
Size: 51
Mode: EHR
Availability: Public∗∗
Size: 105,600
Type: Specific
Use of proprietary model limits evaluation on synthetic patient data.
Rybinski and Karimi (2023); Rybinski et al. (2024)
@@
Jullien et al. (2024)
LLM-guided basic attribute extraction for trial retrieval and filtering. Criterion-level LLM predictions fed to set-reasoning-based re-rank scoring functions.
Availability: Public (TREC CT 2022) LLMs underperform in exclusion criteria labelling.
@@
Availability: Public (Koopman and Zuccon (2016), TREC CT 2021, 2022)
@@
Datta et al. (2025)
Lexical retrievers use LLM generated query expansion. LLM predicts trial-level label and a criterion-level rationale with a matching score. Final ranking based on matching score.
Availability: Public (TREC CT 2023) Trials prefiltered by disorder-specific keywords.
Manually curated disorder-specific instructions.
Model predicted trial-level label is not evaluated
Saeidi et al. (2023); Saeidi (2025) Embed patients and trials to concept vector space. Use variations of Equation 6.1 to compute relevance scores. Availability: Public (TREC CT 2023) No direct connection between criterion-level LLM predictions and trial-level relevance score computations